# xTrimoPGLM: Unified 100B-Scale Pre-trained Transformer for Deciphering the Language of Protein

### v1

@@ Introduction
In this section, we propose xT-Fold, where building on the xTrimoPLGM-100B framework, marks a significant advancement by achieving SOTA results for the PLM-based structure prediction model on benchmarks such as CAMEO and the latest CASP15. Notably, the model offers both high accuracy and computational efficiency, utilizing 4-bit quantization and FlashAttention (?) techniques to run effectively on a single A100 GPU card (SM Section 1.9) This balance of speed and precision makes xT-Fold a tool choice for fast-paced research and drug discovery.
Figure 3: Structure Prediction with xT-Fold. A. xT-Fold’s architecture utilizes a Multi-Layer Perceptron (MLP) connecting language model representations to folding modules, which then generate 3D coordinates and confidence scores (pLDDT). B. TM-score benchmarks for structure prediction models. The bar chart shows the performance of single-sequence PLM-based models and MSA-based models on CAMEO and CASP15 datasets. C. Inference time comparison across models for varying sequence length intervals, showing xT-Fold, ESMFold, OmegaFold, AlphaFold, RoseTTAFold, and MSA search times in seconds.D. Scatter plots compare xT-Fold predictions (x-axis) to other models (y-axis), color-coded by perplexity (green for high, purple for low).
@@
The overall architecture of xT-Fold closely resembles that of ESMFold (Figure 3A). It involves training Multi-Layer Perceptron (MLP) layers to map the respective single representation and pair representation to $d$ and $d^{\prime}$ from $D$ dimension, which are fed into the structure module for 3D coordinates prediction. xT-Fold performs at most three recycling steps before outputting the final atomic-level structure.
@@
We evaluated two individual test sets, CAMEO and CASP15, both of which are out-of-distribution in terms of their timelines, ensuring differences from our training set (SM Section 1.10). CAMEO includes 194 samples (released date between April and June 2022), while CASP15 consists of 56 publicly available proteins (released in May 2022). Initially, we observed the perplexity (the lower are better) of these two datasets on PLMs, which served as the backbones of the folding models (Figure S13). On CAMEO and CASP15, xTrimoPGLM achieved a perplexity of 4.01 and 4.45, respectively, in contrast to ESMFold’s PLM (ESM2-3B), which scored 5.21 and 6.18. The perplexity (PPL) demonstrates that language models generally have a better understanding of the CAMEO dataset compared to CASP15. This suggests that CAMEO might be less challenging than CASP15, indicating that CASP15, relative to CAMEO, is closer to OOD data for language models.
@@
We conducted a comparative analysis against other MSA-free PLM-based models (Figure 3B), including ESMFold, OmegaFold, and 4-bit xT-Fold. In performance evaluation, xT-Fold achieved a TM-score of 0.86 on the CAMEO dataset and 0.70 on CASP15. The scores for ESMFold were 0.85 and 0.65, respectively, while OmegaFold scored 0.80 and 0.60 on these datasets. When compared to methods that integrate MSAs and template retrievals, such as AlphaFold2 and RosettaFold, PLM-based models like xT-Fold and ESMFold showed performance closely matching that of RosettaFold on the CAMEO dataset. However, on the more OOD dataset, i.e. CASP15, the overall efficacy of the PLM-based methods still lagged behind MSA-augmented approaches.
@@
On the other hand, when benchmarked on the CASP15 and CAMEO test sets across various sequence length intervals, the inference speed of MSAs/template-based models is notably slower than that of PLM-based models (as shown in Figure 3C), lagging by approximately 10x to 50x. This disparity remains even for accelerated AlphaFold2, which is optimized with FlashAttention (FA), achieving an acceleration of 2x to 8x compared to the open-source variant (available at https://github.com/google-deepmind/alphafold) across sequence lengths ranging from 200 to 1000. For the PLM-based methods, the open-source versions of ESMFold (https://github.com/facebookresearch/esm) and OmegaFold (https://github.com/HeliXonProtein/OmegaFold) were utilized. As result, xT-Fold overall exhibits a marginal speed advantage over ESMFold and OmegaFold, due to the optimization with FA. Particularly on longer sequence intervals, xT-Fold demonstrates its superiority. The relative slowness of OmegaFold is attributed to its default setting of using 10 recycling passes to achieve better results.
@@
Upon further examination of the scatter plots (Figure 3D), a perceptible correlation between the perplexity from xTrimoPGLM and the structural metric TM-score is observed. Notably, for samples with intermediate PPL values, xT-Fold demonstrates enhanced performance compared to ESMFold and OmegaFold. In the third column of Figure 3D, xT-Fold is juxtaposed with AlphaFold2 with single sequence input. Given that AlphaFold2 is not trained on single sequences, a significant drop in overall effectiveness is evident. The last two columns compare xT-Fold against AlphaFold2 and RosettaFold. It is apparent that where PPL is moderately high, corresponding to less accurate predictions, xT-Fold’s performance is less effective than the two methods, For samples with lower PPL predictions, xT-Fold occasionally surpasses RosettaFold. In summary, while scaling single-sequence models enhances the performance, but still struggles with OOD data, which is effectively addressed by MSAs augmentation.
@@
Autoregressive models have emerged as powerful tools for representing the diverse array of evolutionary sequences found in nature. This capability facilitates the generation of novel protein sequences, exhibiting diverse folding patterns that significantly diverge from naturally occurring proteins (?, ?, ?). To validate the generative ability of xTrimoPGLM-100B, we conduct an extensive analysis of the properties of protein sequences synthesized by xTrimoPGLM-100B under various generative scenarios. Our investigation spans several generative contexts: universal protein synthesis utilizing pre-training data and task-specific sequence generation via Supervised Fine-Tuning (SFT), sequence creation following Reinforcement Self-Training (ReST). This multifaceted approach provides deep insights into the potential and versatility of xTrimoPGLM-100B in advancing protein sequence generation.
Figure 4: Diversification of Generated Proteins by xTrimoPGLM. A. Structural Prediction Confidence and Natural Protein Analogues: We assess the xT-Fold predicted confidence (pLDDT scores), the resemblance to proteins cataloged in the Protein Data Bank (TM-score and sequence identity) of generated sequences by xTrimoPGLM (green, N=2,861) and ProGen2-xlarge (blue, N=2,759). The violin plots mark the median, upper and lower quartiles, and 1.5× interquartile range (whiskers). B. Cartography of Protein Structure Space: We present a comprehensive mapping of protein structural space as informed by sequences generated by xTrimoPGLM. Each node represents a sequence generated by xTrimoPGLM or a sequence from SCOPe70_2.08. Two nodes are linked when one of them can be searched from SCOPe70_2.08 with an alignment of at least 20 amino acids and 70% hhsearch probability. The color coding corresponds to distinct SCOP structural classes, with xTrimoPGLM-generated sequences highlighted in white. For illustrative purposes, we extract and show 6 examples from generated sequences. The the PDB chain ID with the highest structural similarity to the generated sequence, their sequence identity and TM-score are displayed above each example. The color of structure matches to the xT-Fold pLDDT values. The blue color represent high confidence (pLDDT>90). 4nnz_A: Probable zinc protease. 2iah_A: Ferripyoverdine receptor. 6vby_A: Cinnamic acid 4-hydroxylase. 4v37_B: Betaine aldehyde dehydrogenase. 2nuu_E: Ammonia channel. 5msq_A: Carboxylic acid reductase. More visualized cases generated by xTrimoPGLM-100B refer to Figure S10.
@@
xTrimoPGLM Generate Sequences with Diverse Structures. To evaluate the generative capacity of xTrimoPGLM , we generate 2,861 sequences with the xTrimoPGLM-100B model utilizing [gMASK] indicators as the inserting prompt. This process generates new sequences by continuously predicting the next token in an auto-regressive manner until the <eos> token is predicted or the pre-set maximum length is reached. We use nucleus sampling (temperature=0.8, top p=0.8) to increase the generation diversity. For comparison, we also generated 2,759 sequences with the ProGen2-xlarge model with the same hyper-parameters (temperature=0.8, top p=0.8).
@@
We employed xT-Fold to predict the structures of all generated sequences. Our model generated proteins with higher confidence scores than ProGen2-xlarge (median pLDDT scores 85.4 vs 80.8, Figure 4A upper). We then used Foldseek to search for the most structurally similar natural proteins from the PDB database. We measured the structural and sequence similarity using TM-score and sequence identity, respectively. Our model exhibited much higher structural resemblance to PDB entries than ProGen2-xlarge (median TM-score of 0.658 vs 0.560) with very low sequence identity (median sequence identity of 0.117 vs 0.079).
This potentially allows us to access a larger sequence in the protein manifold while we try to design certain protein structures. Yet, the critical question remains: Does xTrimoPGLM have the capability to not only replicate but also expand the known protein universe, introducing novel topologies and functions? Figure 4B showcases a comprehensive network of the protein structural space, informed by sequences synthesized by xTrimoPGLM. Each node corresponds to a xTrimoPGLM-generated sequence or a sequence from SCOPe70_2.08. Sequences originating from xTrimoPGLM are distinctly marked in white. This network vividly illustrates that xTrimoPGLM successfully generates novel protein sequences encompassing a broad spectrum of protein folds, while maintaining low sequence identity, indicating the model’s potential in pioneering new functional and structural realms. A further illustration of additional generative examples and hyper-parameter analysis refers to SM Section 1.12.
@@
Figure 5: Robust Alignment Capabilities of xTrimoPGLM in Protein Sequence Generation towards Desired Properties. Quantitative result in five selective tasks: We present a detailed quantitative analysis of xTrimoPGLM, enhanced through Supervised Fine-Tuning (SFT) and Reinforcement Self-Training (ReST), across five specifically chosen tasks. The analysis highlights xTrimoPGLM’s proficiency in achieving alignment with specific task objectives, as evidenced by the average scores (where higher scores indicate superior alignment performance).
@@
Enhanced Protein Sequence Generation through Supervised Fine-Tuning and Reinforcement Self-Training. The xTrimoPGLM-100B excels in generating diverse protein sequences but faces challenges in aligning to produce sequences with specific properties or families, such as lysozymes or immunoglobulins. This limitation is a critical bottleneck for applications in various industries, including pharmaceuticals and agriculture. Adopting strategies from OpenAI’s GPT models (?), xTrimoPGLM-100B serves as a protein foundational model, equipped with vast knowledge from trillions of residue tokens. We apply established alignment methods like Supervised Fine-Tuning (SFT) (?) on select protein families and enhance this with Reinforcement Self-Training (ReST) (?), based on the SFT model. We fine-tune xTrimoPGLM-100B on datasets representing common protein structures or chemical properties. We choose five tasks from 18 benchmark protein understanding tasks (Fold Prediction, Temperature Stability, Localization Prediction, Fluorescence Prediction, and Fitness Prediction) for fine-tuning, detailed in SM Section 1.11. Comparative analysis is conducted against ProtGPT2 and ProGen2 using identical SFT protocols. To circumvent trivial results, we employ the non-fine-tuned xTrimoPGLM-100B as a control. Task-specific predictors evaluate the quality of the generated sequences, acting as biased evaluators. Due to the impracticality of in vitro validation, we used in silico simulators, a common approach in prior research (?, ?). Our findings, illustrated in Figure 5, reveal that: 1). Sequences from xTrimoPGLM with SFT consistently score higher on targeted properties than the non-fine-tuned baseline, 2). xTrimoPGLM surpasses ProGen2 and ProtGPT2 under the same SFT conditions for most tasks, underscoring its efficacy as a foundational model capable of superior alignment with minimal data or fewer tuning steps, in line with observed scaling behavior. Furthermore, we implement a one-step Reinforcement Self-Training process, which utilizes task predictors as reward models, guiding the self-training of the SFT-enhanced xTrimoPGLM-100B as follows: 1). Task predictors evaluate the quality of sequences from the SFT model, 2). Sequences of higher quality are then selected to form a new dataset, which is used for further fine-tuning. This iterative process results in the development of the Reinforcement Self-Training (ReST) model. Remarkably, the ReST model effectively synthesizes sequences closely resembling natural datasets, highlighting xTrimoPGLM’s potential as a robust protein synthesizer for industrial applications.
@@
To further showcase that xTrimoPGLM can generate proteins that mimic the natural sequences with the SFT alignment pipeline. We fine-tune the model on four SCOP fold-type sequences. Quantitative analyses exploring the relationship between structural prediction confidence, as indicated by xT-Fold’s pLDDT scores, sequence identities, and their structural counterparts in the Protein Data Bank (PDB), as measured by TM-score, are presented in Figure 6. These findings underscore xTrimoPGLM’s exceptional capability in generating protein sequences that not only embody specific structural characteristics but also align closely with established PDB entries, thereby reinforcing its potential as a tool for synthesizing proteins with targeted structural attributes.
@@
Figure 6: Cases Study of xTrimoPGLM in Controllable Generation. Generation of four SCOP fold types by xTrimoPGLM (SFT): Ferredoxin-like (A), Immunoglobulin-like beta-sandwich (B), SH3-like barrel (C) and SDR-type extended Rossmann fold (D). The structures with interleaved green, blue and gray color represents generated proteins. And the structures with gold color are the most structurally similar proteins from PDB database. Percentages in parentheses indicate sequence identity.
Limitations
Challenges of Model Scalability

### v2

@@ Results
In this section, we propose xT-Fold, where building on the xTrimoPLGM-100B framework, marks a significant advancement by achieving SOTA results for the PLM-based structure prediction model on benchmarks such as CAMEO and the latest CASP15. Notably, the model offers both high accuracy and computational efficiency, utilizing 4-bit quantization and FlashAttention  (?) techniques to run effectively on a single A100 GPU card (Supplementary Section 2) This balance of speed and precision makes xT-Fold a tool choice for fast-paced research and drug discovery.
@@
The overall architecture of xT-Fold closely resembles that of ESMFold (Figure 3A). It involves training Multi-Layer Perceptron (MLP) layers to map the respective single representation and pair representation to $d$ and $d^{\prime}$ from $D$ dimension, which are fed into the structure module for 3D coordinates prediction. We used a 48-layer evoformer with approximately 88 million parameters. The structure module accounts for about 2 million parameters, plus additional heads (81K), bringing the total to around 90 million parameters. The maximum recycle count is 3. During training, the recycle count was randomly selected between 0 and 3, consistent with the original AF2 and ESMFold approaches.
@@
We evaluated two individual test sets, CAMEO and CASP15, both of which are out-of-distribution in terms of their timelines, ensuring differences from our training set (Supplementary Section 3). CAMEO includes 194 samples (released date between April and June 2022), while CASP15 consists of 56 publicly available proteins (released in May 2022). Initially, we observed the perplexity (the lower are better) of these two data sets in PLM, which served as the backbone of the folding models (Supplementary Figure 8). On CAMEO and CASP15, xTrimoPGLM achieved a perplexity of 4.01 and 4.45, respectively, in contrast to ESMFold’s PLM (ESM2-3B), which scored 5.21 and 6.18. The perplexity (PPL) demonstrates that language models generally have a better understanding of the CAMEO dataset compared to CASP15. This suggests that CAMEO might be less challenging than CASP15, indicating that CASP15, relative to CAMEO, is closer to OOD data for language models.
@@
We conducted a comparative analysis against other MSA-free PLM-based models including ESMFold, OmegaFold, and 4-bit xT-Fold (Figure 3B and Supplementary Table 1). In performance evaluation, xT-Fold achieved a TM-score of 0.86 on the CAMEO dataset and 0.70 on CASP15. The scores for ESMFold were 0.85 and 0.65, respectively, while OmegaFold scored 0.80 and 0.60 on these datasets. When compared to methods that integrate MSAs and template retrievals, such as AlphaFold2 and RosettaFold, PLM-based models like xT-Fold and ESMFold showed performance closely matching that of RosettaFold on the CAMEO dataset. However, on the more OOD dataset, i.e. CASP15, the overall efficacy of the PLM-based methods still lagged behind MSA-augmented approaches.
@@
On the other hand, when benchmarked on the CASP15 and CAMEO test sets across various sequence length intervals, the inference speed of MSAs/template-based models is notably slower than that of PLM-based models (Figure 3C), lagging by approximately 10x to 50x. This disparity remains even for accelerated AlphaFold2, which is optimized with FlashAttention (FA), achieving an acceleration of 2x to 8x compared to the open-source variant (available at https://github.com/google-deepmind/alphafold) across sequence lengths ranging from 200 to 1000 and kept the output consistent with the canonical self-attention. For the PLM-based methods, the open-source versions of ESMFold (https://github.com/facebookresearch/esm) and OmegaFold (https://github.com/HeliXonProtein/OmegaFold) were utilized. As result, xT-Fold overall exhibits a marginal speed advantage over ESMFold and OmegaFold, due to the optimization with FA. Particularly on longer sequence intervals, xT-Fold demonstrates its superiority. The relative slowness of OmegaFold is attributed to its default setting of using 10 recycling passes to achieve better results.
@@
Upon further examination of the scatter plots (Figure 3D), a perceptible correlation between the perplexity from xTrimoPGLM and the structural metric TM-score is observed. Notably, for samples with intermediate PPL values, xT-Fold demonstrates enhanced performance compared to ESMFold and OmegaFold. xT-Fold is also juxtaposed with AlphaFold2 with single-sequence input. Given that AlphaFold2 is not trained on single sequences, a significant drop in overall effectiveness is evident. The last two columns compare xT-Fold with AlphaFold2 and RosettaFold. It is apparent that where PPL is moderately high, corresponding to less accurate predictions, xT-Fold’s performance is less effective than the two methods, For samples with lower PPL predictions, xT-Fold occasionally surpasses RosettaFold. To provide a more comprehensive and quantified understanding of the relationship between sequence perplexities (PPL) and the corresponding TM-scores predicted by xT-Fold, we have calculated the Pearson correlation coefficient and P-value for the CAMEO, CASP14, and CASP15 datasets. The Pearson coefficients and p-value pair are (-0.415, 1.7e-9), (-0.579, 1e-5), and (-0.239, 8e-2), respectively. These results clearly indicate negative correlations between PPLs and TM-scores on CAMEO and CASP14, but not strictly correlations in CASP15. Overall, these quantified metrics demonstrate that lower perplexities are associated with more accurate predicted structures. In summary, while scaling single-sequence models enhances the performance, it still struggles with OOD data, which is effectively addressed by MSAs augmentation.
Figure 4: Diversification of Generated Proteins by xTrimoPGLM. A. Violin plots comparing ESMFold-predicted confidence (pLDDT scores) and similarity to Protein Data Bank (PDB) entries—measured by TM-score and sequence identity—for sequences generated by xTrimoPGLM (green, $N$ = 14,626) and PROGEN2-xlarge (blue, $N$ = 8,466). The plots display the median, upper and lower quartiles, and whiskers representing 1.5$\times$ the interquartile range. B. The comprehensive mapping of protein structural space as informed by sequences generated by xTrimoPGLM. Each node represents a sequence generated by xTrimoPGLM or a sequence from SCOPe70_2.08. Two nodes are linked when one of them can be searched from SCOPe70_2.08 with an alignment of at least 20 amino acids and 70% hhsearch probability. The color coding corresponds to distinct SCOP structural classes, with xTrimoPGLM-generated sequences highlighted in white. For illustrations (Supplementary Figure 11), we showcase 6 examples from generated sequences. The PDB chain ID with the highest structural similarity to the generated sequence, their sequence identity, and TM-score are displayed above each example. The color of the structure matches the xT-Fold pLDDT values. The blue color represents high confidence (pLDDT>90). 4nnz_A: Probable zinc protease. 2iah_A: Ferripyoverdine receptor. 6vby_A: Cinnamic acid 4-hydroxylase. 4v37_B: Betaine aldehyde dehydrogenase. 2nuu_E: Ammonia channel. 5msq_A: Carboxylic acid reductase.
@@
Autoregressive models have emerged as powerful tools for representing the diverse array of evolutionary sequences found in nature. This capability facilitates the generation of novel protein sequences, exhibiting diverse folding patterns that significantly diverge from naturally occurring proteins  (?, ?). To validate the generative ability of xTrimoPGLM-100B, we conduct an extensive analysis of the properties of protein sequences synthesized by xTrimoPGLM-100B under various generative scenarios. Our investigation spans several generative contexts: universal protein synthesis utilizing pre-training data and task-specific sequence generation via Supervised Fine-Tuning (SFT), sequence creation following Reinforcement Self-Training (ReST). This multifaceted approach provides deep insights into the potential and versatility of xTrimoPGLM-100B in advancing protein sequence generation.
@@
xTrimoPGLM Generate Sequences with Diverse Structures. To evaluate the generative capacity of xTrimoPGLM , we generate 14,626 sequences with the xTrimoPGLM-100B model utilizing [gMASK] indicators as the inserting prompt. This process generates new sequences by continuously predicting the next token in an auto-regressive manner until the <eos> token is predicted or the pre-set maximum length is reached. We use nucleus sampling by combining different top P (0.5, 0.7, 0.9, 1.0) and sampling temperature (0.2, 0.4, 0.6, 0.8, 1.0) parameters. For each parameter combination, we generated 2,000 protein sequences and limited the maximum length to 800 tokens. Then we performed a simple filtering on the generated sequences: (1) removing the generated sequences with perplexity>10; (2) removing the sequences containing repeated fragments; (3) MMseqs2 clustering (–min-seq-id 0.9 -c 0.5) and only leave the centroid sequence. We also used same strategy to generate 8,466 sequences for PROGEN2.
@@
We employed ESMFold to predict the structures of all generated sequences. Our model generated proteins with higher confidence scores than PROGEN2-xlarge (mean pLDDT scores 84.0 vs 74.3, Figure 4A upper). We then used Foldseek to search for the most structurally similar natural proteins from the PDB database. We measured the structural and sequence similarity using TM-score and sequence identity, respectively. Our model exhibited much higher structural resemblance to PDB entries than PROGEN2-xlarge (mean TM-score 0.695 vs 0.522) with very low sequence identity (mean sequence identity 0.224 vs 0.165) and high diversity (Supplementary Figure 9A). We have stratified the generated sequences into four groups based on perplexity values (<2, 2-5, 5-8, >8). Across all perplexity ranges, our model consistently produces proteins with higher confidence scores and greater structural and sequence resemblance to PDB entries compared to PROGEN2 (Supplementary Figure 10).
@@
Figure 5: Robust Alignment Capabilities of xTrimoPGLM in Protein Sequence Generation towards Desired Properties. Quantitative analysis of xTrimoPGLM, enhanced with Supervised Fine-Tuning (SFT) and Reinforcement Self-Training (ReST), across five selected tasks. The number of sampled generated sequences, $N$(other), and natural sequences, $N$(nature), used in the analysis are illustrated in the figure. The results demonstrate xTrimoPGLM’s effectiveness in aligning with specific task objectives, as shown by the average scores (higher scores indicate better alignment). P.Gen2 refers to the PROGEN2-xlarge model  (?) with 6.4 billion parameters, and P.GPT2 denotes the ProtGPT2 model  (?) with 740 million parameters.
To measure the generated protein distribution in the protein space, we used Foldseek to align the structures predicted by ESMFold to the AlphaFold/UniProt50-minimal structure dataset to obtain the maximum TM-score for each generated sequence. This dataset is constructed by MMseqs2 with 50% sequence similarity from the UniProt subset in the AlphaFold Database. To obtain the maximum sequence similarity to natural protein sequences, we also retrieved each generated sequence from UniProt, UniClust30, and BFD database with MMseqs2 and HHblits (Supplementary Figure 9B). The UniProt and BFD databases contain more than 2.5 billion protein sequences, representing the space of currently known protein sequences in nature. For sequences with ESMFold pLDDT>80 (N=11,048), we show the maximum sequence similarity and maximum TM-score of all query sequences in the scatter plot (Supplementary Figure 9C), with a mean sequence similarity value of 0.699 and a mean TM-score value of 0.864. The vast majority of generated structures have similar 3D structures to those in UniProt50, which confirms that our pre-trained model comprehensively understands and represents the protein universe.
This potentially allows us to access a larger sequence in the protein manifold while we try to design certain protein structures. Figure 4B showcases a comprehensive network of the protein structural space, informed by sequences synthesized by xTrimoPGLM. Each node corresponds to a xTrimoPGLM-generated sequence or a sequence from SCOPe70_2.08. Sequences originating from xTrimoPGLM are distinctly marked in white. This network vividly illustrates that xTrimoPGLM successfully generates novel protein sequences encompassing a broad spectrum of protein folds, while maintaining low sequence identity. More analysis refers to Supplementary Section 6.
@@
Enhanced Protein Sequence Generation through Supervised Fine-Tuning and Reinforcement Self-Training. The xTrimoPGLM-100B excels in generating diverse protein sequences but faces challenges in aligning to produce sequences with specific properties or families, such as lysozymes or immunoglobulins. This limitation is a critical bottleneck for applications in various industries, including pharmaceuticals and agriculture. Adopting strategies from OpenAI’s GPT models  (?), xTrimoPGLM-100B serves as a protein foundational model, equipped with vast knowledge from trillions of residue tokens. We apply established alignment methods like Supervised Fine-Tuning (SFT)  (?) on select protein families and enhance this with Reinforcement Self-Training (ReST)  (?), based on the SFT model. We fine-tune xTrimoPGLM-100B on datasets representing common protein structures or chemical properties. We choose five tasks from 18 benchmark protein understanding tasks (Fold Prediction, Temperature Stability, Localization Prediction, Fluorescence Prediction, and Fitness Prediction) for fine-tuning. Specifically, we adopted two filtering strategies to obtain the SFT datasets: Regression Tasks (Fluorescence and Fitness Prediction): We filtered samples whose label scores exceed a certain threshold, then fine-tuned the models with these samples. Classification Tasks (Fold Prediction, Temperature Stability, Localization): We used samples from one category to fine-tune the model and generate new samples. Moreover, we further filter the protein sequences generated by the SFT models as the ReST datasets. We applied the same filtering and fine-tuning settings during the Reinforcement Self-Training (ReST) stages. We finetune our model and the baseline models, such as PROGEN2 and ProtGPT2, using the same causal language modeling regime (Supplementary Section 5). Comparative analysis is conducted against ProtGPT2 and PROGEN2 using identical SFT protocols. To circumvent trivial results, we employ the non-fine-tuned xTrimoPGLM-100B as a control. Task-specific predictors evaluate the quality of the generated sequences, acting as biased evaluators. Due to the impracticality of in vitro validation, we used in silico simulators, a common approach in prior research  (?, ?). More specifically, we use the corresponding task predictor to predict the scores of the desired class (for classification tasks) or the regression scores (for regression tasks) to validate whether the SFT models could generate sequences with the desired properties.
Our findings reveal that: 1). Sequences from xTrimoPGLM with SFT consistently score higher on targeted properties than the non-fine-tuned baseline, 2). xTrimoPGLM surpasses PROGEN2 and ProtGPT2 under the same SFT conditions for most tasks, underscoring its efficacy as a foundational model capable of superior alignment with minimal data or fewer tuning steps, in line with observed scaling behavior (Figure 5). Furthermore, we implement a one-step Reinforcement Self-Training process, which utilizes task predictors as reward models, guiding the self-training of the SFT-enhanced xTrimoPGLM-100B as follows: 1). Task predictors evaluate the quality of sequences from the SFT model, 2). Sequences of higher quality are then selected to form a new dataset, which is used for further fine-tuning. This iterative process results in the development of the Reinforcement Self-Training (ReST) model. Remarkably, the ReST model effectively synthesizes sequences closely resembling natural datasets, highlighting xTrimoPGLM’s potential as a robust protein synthesizer for industrial applications.
@@
To further showcase that xTrimoPGLM can generate proteins that mimic the natural sequences with the SFT alignment pipeline. We fine-tune the model on four SCOP fold-type sequences. Quantitative analyses exploring the relationship between structural prediction confidence, as indicated by xT-Fold’s pLDDT scores, sequence identities, and their structural counterparts in the Protein Data Bank (PDB), as measured by TM-score (Figure 6). These findings underscore xTrimoPGLM’s exceptional capability in generating protein sequences that not only embody specific structural characteristics but also align closely with established PDB entries, thereby reinforcing its potential as a tool for synthesizing proteins with targeted structural attributes.
@@
Figure 6: Cases Study of xTrimoPGLM in Controllable Generation. Generation of four SCOP fold types by xTrimoPGLM (SFT): Ferredoxin-like (A), Immunoglobulin-like beta-sandwich (B), SH3-like barrel (C), and SDR-type extended Rossmann fold (D). Generated protein structures are depicted in interleaved green, blue, and gray colors, whereas gold-colored structures represent the most structurally similar proteins from the PDB database. Percentages in parentheses indicate sequence identity.
Discussion